OpenStudy (anonymous):

Can you explain how to do this? (2x^2y^3)(-3x^3y^4)(6xy)=?

OpenStudy (amistre64):

id stack them to keep track of it better

OpenStudy (anonymous):

What do you mean?

OpenStudy (amistre64):

2 x^2 y^3 -3 x^3 y^4 6 x y i mean, id stack them so that i could keep track of them better :/

OpenStudy (anonymous):

ohh okay

OpenStudy (jiteshmeghwal9):

\[(2.-3.6)x^2y^3x^3y^4xy\]

OpenStudy (anonymous):

my bad

OpenStudy (amistre64):

like bases add exponents; and constants just get mulitplied together

OpenStudy (jiteshmeghwal9):

multiply the constants & multiply the like terms to gt ur answer

OpenStudy (jiteshmeghwal9):

gt it @maherman97 ???

OpenStudy (anonymous):

no... haha what do you mean the constants?

OpenStudy (jiteshmeghwal9):

constants=numerical terms

OpenStudy (jiteshmeghwal9):

such as 3,5 etc.

OpenStudy (anonymous):

well can you walk me through the steps?

OpenStudy (anonymous):

i would start by multiplying (6xy) by 2x^2y^3) get the answer and the multiplying the answer by (-3x^3y^4)

OpenStudy (anonymous):

how do you multiply the first part?

OpenStudy (anonymous):

ok . (6xy)(2x^2y^3)= 12x^3y^4 using exponents rules.

OpenStudy (anonymous):

okay so then why do i multiply that by (-3x^3y^4)

OpenStudy (anonymous):

that what the operator dose.

OpenStudy (anonymous):

the what???

OpenStudy (anonymous):

im soooo confused!

OpenStudy (anonymous):

(A)(B)(C) means A multiply B multiply C as an example .

OpenStudy (anonymous):

rightt....

OpenStudy (anonymous):

So then i multiply 12x^3y^4 by (-3x^3y^4)

OpenStudy (anonymous):

SO THAS why we use the operator (multiplication) thats what i mean by "operator" hope that helps

OpenStudy (anonymous):

okay so can you show me the steps into multiplying 12x^3y^4 by (-3x^3y^4)

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

so what do i do first?

OpenStudy (anonymous):

1>multiply 12by -3 = -36 2> x^3by x^3 = x^6 3> y^4 by y^4= y^8

OpenStudy (anonymous):

so it would be -36x^8y^8?

OpenStudy (anonymous):

x^6* my badd

OpenStudy (anonymous):

okay thanks soooo much!

OpenStudy (anonymous):

ive got another for you

OpenStudy (anonymous):

i would appreciate a "BEST REPONSE " if you think so?? :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sooo if i have \[(3x)^{3}(4x)^{2}\] would the answer be 216?

OpenStudy (anonymous):

no (3x)3(4x)2 thats not the same. be careful mate.. :)

OpenStudy (anonymous):

so what would it be? and what did i do wrong?

OpenStudy (anonymous):

\[(3x)^3\] is not the same as \[3x^2\] = henc \[(3x)^3\]=\[3^3 x^3= 27x^3\] and \[3x^3=3x^3\]